The Sum Of ages of a father and his son is 60.
2 years ago the age of the father was three times as old as his son.
what are their ages?

Answer by Cromlix(4381)

Hi there,
The Sum Of ages of a father and his son is 60.
F + S = 60
F = 60 - S
2 years ago the age of the father
was three times as old as his son.
F - 2 = 3(S - 2)
Replace F with (60 - S)
(60 - S) - 2 = 3(S - 2)
60 - S - 2 = 3S - 6
Collect like terms:
- S - 3S = - 6 + 2 - 60
- 4S = - 64
4S = 64
S = 16
Son's age is 16 years old
Father's age = 44 years old
Hope this helps:-)